Two particles A and B are projrcted from the same point on the ground. Particle A is projecyed on a smooth horizontal surface with speed v, whereas B is projrcted with velocity 2v/√3. The two particles collide when the particle B drops on the surface for the first time.
Find:
a) The angle with which the particle B was projected into air.
b) The separation between the two particles when B is at its highest position.

Grade:11

1 Answers

Eshan
5 years ago
Dear student,

Since the particle collide on the horizontal surface after finite time, their horizontal displacement must be same. Hence the horizontal component of their velocities must be same. Therefore\dfrac{2v}{\sqrt{3}}cos\theta=v\implies cos\theta=\dfrac{\sqrt{3}}{2}\implies \theta=30^{\circ}
When it is at maximum height, the separation is the height of the particle B which is equal to\dfrac{u^2sin^2{\theta}}{2g}=\dfrac{v^2}{6g}